# AI-Driven Predictive Analytics Approach for Early Prognosis of Chronic Kidney Disease Using Ensemble Learning and Explainable AI thanks:

### v1

@@ III-C Statistical Measures on Feature Significance
[…]
TABLE V: Statistically Significant Features as Found by the Logit Model
[…]
@@
TABLE VI: Selected Set of Features from Filter-based and Wrapper Methods
@@
Information Gain ‘pcv’, ‘hemo’, ‘rbcc’, ‘sg’, ‘sc’, ‘al’, ‘htn’, ‘sod’, ‘rbc’, ‘dm’, ‘pot’
Variance Threshold ‘sg’, ‘pcc’, ‘ba’, ‘htn’, ‘cad’, ‘appet’, ‘pe’, ‘ane’
Forward Selection ‘bp’, ‘sg’, ‘al’, ‘rbc’, ‘bgr’, ‘sc’, ‘pot’, ‘hemo’, ‘wbcc’, ‘rbcc’, ‘ane’
Recursive Feature Elimination ‘al’, ‘rbc’, ‘pc’, ‘ba’, ‘sc’, ‘pot’, ‘hemo’, ‘htn’, ‘cad’, ‘appet’, ‘pe’
@@ III-E Feature Selection Process
Since 25 features were initially present for the dataset, it is important to identify features which provide little information in terms of variability in the prediction. At first, several filter based methods of feature selection were applied, then the wrapper methods for the feature selection were employed. In this way, it was easier for us to determine which features are contributing more in the prediction of the CKD.
TABLE VII: Accuracy Metrics of Applied Machine Learning Models
@@
LR 0.93 1.0 0.96 0.97
NV 0.95 0.98 0.96 0.97
SVM 0.95 1.0 0.97 0.98
Adaboost 0.95 1.0 0.97 0.95
XGBoost 0.97 1.0 0.98 0.99
DT 0.91 1.0 0.95 0.96
RF 0.97 1.0 0.98 0.98
TABLE VIII: Accuracy Metrics of Applied Machine Learning Models with Selected Features
Model Precision Recall F-1 Score Accuracy
LR 0.97 0.97 0.97 0.98
NV 0.97 1.0 0.98 0.99
SVM 0.97 0.97 0.97 0.98
Adaboost 0.97 0.97 0.97 0.98
XGBoost 0.97 1.0 0.98 0.99
DT 0.86 0.94 0.90 0.93
RF 0.98 1.0 0.99 0.99
@@
The information gain method was first employed which calculates reduction of entropy or which features retain more information in prediction of the target feature. Then the Variance Threshold method where a threshold of 0.25 was selected as the criteria for feature elimination in order to remove constant and quassi-constant features. These features have similar values across all observations in the dataset and provide little to no information in the prediction of a target feature. Both the forward and backward feature elimination yielded same set of variables so only the forward selection results were presented. Finally, the Recursive Feature Elimination (RFE) was used to prune the least important set of features recursively from the current set of features.
@@
From the feature selection process, 14 features were narrowed down as found from the Co-relation matrix and the filter, wrapper methods of feature selection. For our model implementation, first some of the traditional Machine Learning models were looked at. Since it’s a binary classification problem, this study would start with Logistic Regression and shift towards using probabilistic models and towards more complicated model like Support Vector Machine. Next, for our explainable AI approach, ensemble tree models like XGradientBoosting, Decision Trees and Random Forests would be explored. The models would be fitted on the selected 14 set of features and then it would be observed how the models are doing in the prediction task.
III-F Discussion with a Nephrologist
@@
After completing the preliminary analysis with the features, it was imperative to discuss with clinical professionals to validate the study or gain research directions. To this end, a visit was paid to Dr. Nazneen Mahmood (MBBS, MD in Nephrology) in Anwar Khan Modern Hospital, Dhaka, Bangladesh to validate whether which of these features are clinically relevant to distinguish CKD vs. Non CKD individuals. As per her findings, Specific Gravity (Sg) and Bacteria (Ba) have very little clinical significance in prediction of CKD. In stead of these features, our access to some features like Glomerular Filtration Rate (eGFR), Bi-Carbonate (HCO3), HbA1C (Mean Blood Glucose), Pro-BNP (B-type natriuretic peptide (BNP) test), S.Ph (Sodium Phosphate), PTH (Parathyroid Hormone) etc. would be more clinically significant variables. So, efforts were established for contacting the Apollo Hospital in Madurai, India several times, specially with the contributing doctor of the dataset but to no avail. So, the scope of this research, is limited to the choice of features by choosing the ones from the feature extraction techniques discussed before and discarding the ones suggested by Dr. Mahmood. So, the final 14 features for applying the Machine Learning models are:‘hemo’, ‘rbcc’, ‘sc’, ‘al’, ‘htn’,‘pot’, ‘rbc’, ‘cad’, ‘appet’, ‘pe’, ‘ane’, ‘pcv’, ‘rbcc’, ‘bgr’,‘wbcc’, ‘class’.
@@ IV Model Application Process
After the exploratory data analysis is done for the dataset, the Machine Learning models were applied. Primarily, this research focused on interpretability of the applied models since this is still a relatively unexplored area of research with chronic kidney diseases. Since the scope of this research lies in just the prediction of CKD addressing the other independent variables, the Machine Learning problem is formulated as a binary classification task. To this end, the Bayesian Estimation i.e. Naive Bayes and Logistic Regression were applied. The Logistic Regression was applied with L1 penalty and ‘Balanced’ class weighted. The ‘Balanced’ class was used to account for the class imbalance as there were only 250 instances of ‘CKD’ and 150 instances of ‘Non-CKD’ in the original dataset and this class imbalance would be increased when the train-test split would be performed for the model applications. The best set of parameters for the Support Vector Machine (SVM) were chosen by random trial basis. The ‘rbf’ kernel with ‘Polynomial’ degree as the parameters produced the best results in the accuracy metrics. Since the scope of this research is focused towards interpretation of the results of the Machine Learning models, the tree-based models from the ensemble family were chosen to apply on the CKD dataset. From the ensemble family, Adaboost Classifier, XGradientBoosting, Decision Tree and Random Forest were chosen. The Decision Trees were applied with the entropy measure by Gini Index. The visualization of the applied Decision Trees provide a good indicator on which features are more important than others in evaluation. The example decision trees shown by Random Forest in figure 2 and figure 3 gives in detail about the Decision Tree branching by Gini Index and also the contribution score of those features for each class of CKD prediction.
@@ V Model Performance on CKD Dataset
Table VII and VIII indicate the scores of the applied models in the accuracy metrics. From the previous sections on feature selection through the 4 techniques in table VI and discussion with expert Nephrologists, final 14 features were narrowed down to compare the prediction performance of the Machine Learning models in the whole feature set versus the selected subset of features. The models discussed in the model application process are evaluated on the whole dataset and also on the selected 14 features.
Figure 2: Confusion Matrix Random Forest
As evident from table VIII, all the applied models have very high scores on the accuracy metrics and the recall score is 1.0 in all the models except LR. It is due to the fact that all models have not predicted any false negatives. So, no patient who has CKD has been misclassified as Non-CKD patient. This is a very good prediction result on this sensitive dataset of CKD prediction. Moreover, the number of false positives are also very low accounting for the high Precision score in all of the models. Overall, all the models have performed really well on the selected subset of 14 features for this dataset. Also, when compared with the Machine Learning models with all 24 features, the subset of 14 features has performed better in all the metrics for all the models. This shows to indicate that the feature engineering process has yielded better selection of features than all combined 24 independent features. Among all the models that are applied on both the entire dataset and selected subset of the features, Random Forest model has performed the best compared to all other models.
@@ V-A Random Forest Model Performance
Since of all the 6 models applied, Random Forest (RF) has performed the best in terms of accuracy metrics, this study explored more of this model performance in terms of confusion matrix, decision trees and then to the explainability.
Looking at the confusion matrix for the Random Forest in figure 2, it is observed that 65 cases of True Negative and 34 cases of True Positive prediction was found in the prediction set. In the contrary, 1 mis-classification occurred with False Positive and no case of False Negative was found. That is why, the precision value is at 0.98 and with no false negative prediction, the recall value is 1.
Figure 3 represents the Decision Trees generated for the Random Forests for the prediction model. The samples at the root or leaf nodes represent the percentage of samples that are reaching that node. The two values in the array represent the contribution of that feature in the prediction of Non-CKD and CKD. The root node at the top dignifies the feature value with the highest Gini Score and the next nodes at the leaves are the feature scores in terms of Gini Impurity in descending order of sample proportion.
Figure 3: Example Decision Tree for Random Forest
It is evident from the decision tree used by the random forest model that it chooses ‘rbcc’, ‘sc’, ‘al’, ‘hemo’, ‘sc’, ‘pcv’ variables in most cases. Point to note here is that Random Forest model combines a number of Decision Trees and based on the best pruning results of the decision trees, the final results are presented in random forest. For this study, the Grid Search technique of the Random Forest model was also employed to find out best set of parameters for prediction. It was found that a depth of 14 with number of estimators as 472, the best Random Forest Model was selected. The best performing Random Forest model would be further evaluated using the feature importances in the training phase and the validation set would be explained using XAI tools.
During the training phase as depicted in figure 4, the Random Forest model has identified the “Hemo”, “Pcv”, “Sc”, “Al” as the features with highest weights. The weights and coefficients of these features from the test and validation set with the help of XAI techniques were analyzed.
@@
In figure 5, a random prediction case was taken where the Random Forest model predicted with 2% confidence that the patient has CKD and with 98% confidence that the patient does not have CKD. The strongly contributing features for the CKD confidence are ‘hemo’, ‘rbc’, ‘al’, ‘htn’, ‘pcv’, ‘pe’ etc. The range of values these features exhibit were also observed when the RF model predicts as Non-CKD with 98% confidence.
For the 2% confidence in CKD prediction, LIME shows only ‘rbcc’ as the contributing feature. Here, the LIME tool has visualized that for this instance of the patient, the best performing Random Forest model has identified it to be a Non-CKD patient because ‘Hemoglobin’ count in gms is greater than 14.5, Red Blood Cell levels are abnormal, low level of Albumin, less chances of Hypertension, Serum Creatinine in a range from 0.9 to 1.30.
@@
If the attention is just focused on few features like “Hemo”, “Sc”, “Al” it can be inferred that the best RF model is confident about non-CKD prediction when the “Hemo” values are greater than 14.5. But it is more confident in CKD prediciton when the values are in a range from 10.9 to 12.5.
@@
Similarly, model is more confident about non-CKD prediction when “Sc” values are are in a range between 0.9 to 1.3 but more confident about CKD prediction when “Sc” values are greater than 1.30 and less than 2.9. Lastly, the feature “Al” values are on the lower end for the non-CKD prediction and on the higher end for the CKD prediction as it should be as per clinical requirements.
From figure 6, LIME has identified that the RF model has 95% confidence that the patient has CKD versus 5% confidence that he does not have CKD. In this case also, there are more features that contribute towards prediction of CKD than prediction of non-CKD. Contrary to figure 6, here the more strongly contributing feature towards positive CKD prediction is ‘rbc’ with a score of 0.14 where it shows that Rbc level is “Abnormal” for that patient. For this patient, ‘Haemoglobin’ count is greater than 10.9 and less than 12.5. Albumin is higher than previous example. Unlike Non-CKD prediction in the previous case, ‘Blood Glucose Random (BGR)’ is a more contributing feature for ‘CKD’ prediction than ‘Non-CKD’ prediction.
In fig 7, f(x) denotes the predicted log odds of CKD patients against E(f(x)) where the SHAP values for individual features give the difference between the predicted log odds and avg. predicted log odds. In log odds measurement, the features ‘hemo’, ‘rbc’, ‘bgr’, ‘sc’, ‘rbcc’ are the features that contributed more for non-CKD prediction denoted by the blue bars. The top contributing features for the Non-CKD prediction are ‘hemo’, ‘rbc’, ‘bgr’, ‘sc’.
TABLE IX: Feature Values and their Contributions from Eli5 for Non-CKD prediction
Feature Name Feature Value Contribution
Hemo 16.5 0.16
Pcv 43.0 0.13
Rbc 2.00 0.104
Al 0.0 0.076
Sc 1.2 0.070
Bgr 125.0 0.042
Htn 1.00 0.020
TABLE X: Feature Values and their Contributions from Eli5 for CKD prediction
Feature Name Feature Value Contribution
Bgr 341.0 0.111
Hemo 12.3 0.104
Sc 1.5 0.099
Rbc 0.0 0.082
Htn 2.0 0.064
Al 1.02 0.056
Pot 4.63 0.008
Fig 8 is designed similar to Figure 7 where in figure 7, the prediction impact of features are shown against their actual values. The top two features are same for both CKD and Non-CKD prediction. The difference of CKD prediction scores to the Non-CKD prediction is that the ‘Presence of Albumin (Al)’ is not providing more impact for Non-CKD prediction where it had more impact on the CKD prediction. The top features are more or less same for both the class prediction but they naturally vary with the actual values of the features.
In table IX and X, the model features with their values and contribution scores are presented for CKD vs. Non-CKD prediction using the Explain to me Like I’m a 5 year old (Eli5) model. The feature values are contribution scores are validated from the earlier implementations of SHAP and LIME plots.
Fig 9 depicts the scatter plots of a test case to visualize the inter-dependence of independent features. Here, the top two features with respect to feature contribution i.e. “rbcc” and “hemo” are analyzed. It is evident from the figure that “rbcc” exhibits a value range from 3.5 to 4.5 when “Hemo” quantity is in a range from 8 to 12. That is a cluster could be imagined within these value ranges for these two features. When the “Hemo” quantity is more towards the higher end starting greater than 12, that’s when another cluster could be imagined with “rbcc” where it can also be seen that “rbcc” values are greater than 5.5 and more towards the higher end for this feature.
Figure 8: Shap Waterfall Plot for Visualizing Global Test Cases
@@
Figure 10 denote the decision making process of the Random Forest model using SHAP in global test cases. The Mean SHAP Value or the Average Impact on model’s output magnitude is shown for both Non-CKD (Denoted by Blue Bars and CKD by red bars). From the validation set after fitting the Random Forest model with the best set of parameters estimated from Grid Search, the features like “Hemo”, “Pcv”, “Rbc”, “Al”, “Sc” are the strongest contributing ones.
The peak values of prediction are also validated by the ALE plots in fig 11. In the X-axis, the effect of prediction is the highest when the ‘Rbcc’ value is at 2 and it plateaus out for the higher values. For Albumin, when the Albumin quantity is at the lowest i.e. at 0, this feature is contributing most towards the prediction and it becomes constant with the increasing values as shown in figure 11.
@@
Figure 11: ALE Plot showing effect on prediction using Al
The Partial Dependence Plots (PDP) in figure 12 validate how the feature values dependency works in the predictions. If the partial dependence plot in fig 12 is analyzed, it would be seen that the more strongly contributing features ‘rbcc’ and ‘hemo’ have their peak contribution when ‘rbcc’ values are greater than 5 and ‘hemo’ values are greater than 15. ‘Hemo’ values also reach a local maxima at values 12.5 to 13. The combined dependence of the two features is also shown when local maxima is reached at 0.41 when the ‘Hemo’ value is at 13, ‘Rbcc’ at 6 and another at 0.55 when the ‘Rbcc’ is at 6 and ‘Hemo’ value greater than 15.

### v2

@@ III-C Statistical Measures on Feature Significance
From table IV, Only those features are deemed as statistically significant in CKD prediction where the ‘P values’ are less the Fisher threshold of 0.005. The same process of “Null Hypothesis” testing is also applied here only in this case the Null Hypothesis refers to the hypothesis that features are strongly significant in the prediction of the target “CKD” feature. Or in other words, the relationship between each of the feature with the target feature is not due to random chance. Looking at the ‘P values’ in table IV, it is evident that all the 7 features denoted here are indicative of the acceptance of the “Null Hypothesis” and rejecting any alternative hypothesis.
TABLE VI: Statistically Significant Features as Found by the Logit Model
Feature Name t-Value P values
@@
TABLE VII: Feature Entropy and Information Gain Values
Feature Name Information Gain Entropy
Hemo 0.45 7.01
Sg 0.43 3.01
Pcv 0.42 5.84
Rbcc 0.42 6.72
Sc 0.42 5.56
Al 0.37 2.80
Sod 0.28 5.81
Rbc 0.26 0.99
Htn 0.23 0.99
Pot 0.21 5.80
TABLE VIII: Selected Set of Features from Filter-based and Wrapper Methods
@@
Information Gain ‘hemo’, ‘pcv’, ‘sg’, ‘rbcc’, ‘sc’, ‘al’, ‘sod’, ‘rbc’, ‘htn’, ‘pot’
Variance Threshold ‘age’, ‘bp’, ‘al’, ‘su’, ‘bgr’, ‘bu’, ‘sc’, ‘sod’, ‘pot’, ‘hemo’, ‘pcv’, ‘wbcc’, ‘rbcc’
Forward Selection ‘age’, ‘bp’, ‘sg’, ‘al’, ‘rbc’, ‘pcc’, ‘ba’, ‘pot’, ‘hemo’, ‘htn’
Recursive Feature Elimination ‘al’, ‘su’, ‘rbc’, ‘sc’, ‘pot’, ‘hemo’, ‘rbcc’, ‘htn’, ‘dm’, ‘pe’
@@ III-E Feature Selection Process
Since 25 features were initially present for the dataset, it is important to identify features would provide little information in terms of variability in the prediction. At first, several filter based methods of feature selection were applied, then the wrapper methods for the feature selection were employed. In this way, it was easier for us to determine which features are contributing more in the prediction of the CKD.
TABLE IX: Accuracy Metrics of Applied Machine Learning Models with Selected Features
@@
LR 0.97 0.97 0.97 0.97
NV 0.95 0.95 0.95 0.94
SVM 0.98 0.98 0.98 0.98
Adaboost 0.97 0.97 0.97 0.97
XGBoost 0.98 0.98 0.97 0.98
DT 0.98 0.98 0.98 0.96
RF 0.98 0.99 0.98 0.98
@@
The information gain method was first employed which calculates reduction of entropy or which features retain more information in prediction of the target feature. The higher the value of entropy, higher is the variance within that feature. So, with information gain we are selecting only those features which have large to moderate entropy and also contribute towards prediction or reducing the determining the uncertainty of target feature as shown in table 5. Then the Variance Threshold method was applied to remove the constant and quassi-constant features. Constant features are the ones which have variance across the entire dataset. We can also say, they are homogenous features. Quassi-constant features are the ones where one value dominates 99% of the records. We need to remove these constant and quassi-constant features because no variance in the feature set would only increase computational complexity of the models, not adding any information in the features. A high threshold of 0.75 was selected as the criteria to remove any feature that has less than 0.75 as the variance. Both the forward and backward feature elimination yielded same set of variables, so only the forward selection results were presented. Logistic Regression (LR) is applied to the entire dataset to see how well a linear model has fits the CKD dataset. So, the process of adding features is stopped when addition of a new feature does not improve the LR fit on the dataset. That is how the forward selection method works. Finally, the Recursive Feature Elimination (RFE) was used to prune the least important set of features recursively from the current set of features. Similar to the forward selection, RFE also uses logistic regression to recursively eliminate features based on feature ranking.
We have implemented two different categories of feature selection process. The first category relies on the variance within features. The two methods in this category are information gain and variance threshold. The common subset of features selected from these two methods are ‘hemo’, ‘pcv’, ‘sc’, ‘al’, ‘sod’, ‘pot’.
Next category of feature selection is through the use Logistic Regression model to find out which features add value to the prediction of target class ”ckd” based on feature scores and ranking. Forward selection adds features if the LR model’s fit improves by addition of features and RFE removes the features based on feature ranking. The common subset of features selected from these two methods are ‘al’, ‘rbc’, ‘pot’, ‘hemo’, ‘htn’.
Since the two categories of feature selection methods are different, hence there are differences in the selected subset of features. Combining the feature correlation, statistical significance and the different methods of feature elimination, we have 6 methods to choose the final subset of features. We would analyze if any feature falls in at least two different methods, so we would include that feature in the final list.
@@
From the feature selection process, 13 features (‘hemo’, ‘sg’, ‘rbc’, ‘al’, ‘htn’, ‘pot’, ‘dm’, ‘pcv’, ‘sc’, ‘rbcc’, ‘sod’, ‘age’, ‘bp’) were narrowed down as found from the Co-relation matrix and the filter, wrapper methods of feature selection. We would discard ‘Sg’, ‘Pcv’, ‘Sod’, ‘Pot’, ‘rbc’, ‘Rbcc’ from the selected subset of features based on our discussions with nephrologists which is analyzed in more detail in the ”Consultation with First Nephrologist” section. For our model implementation, first some of the traditional Machine Learning models were looked at. Since it’s a binary classification problem, this study would start with Logistic Regression and shift towards using probabilistic models and towards more complicated model like Support Vector Machine (SVM). Next, for our explainable AI approach, ensemble tree models like XGradientBoosting (XGBoost), Decision Trees and Random Forests would be explored. The models would be fitted on the selected 6 features and then it would be observed how the models are doing in the prediction task.
III-F Consultation with First Nephrologist
@@
After completing the preliminary analysis with the features, it was imperative to discuss with clinical professionals to validate the study or gain research directions. To this end, a visit was paid to Dr. Nazneen Mahmood (MBBS, MD in Nephrology) in Anwar Khan Modern Hospital, Dhaka, Bangladesh to validate whether which of these features are clinically relevant to distinguish CKD vs. Non CKD individuals. As per her findings, Specific Gravity (Sg) and Bacteria (Ba) have very little clinical significance in prediction of CKD. Moreover, the features indicating serum electrolytes like ‘Sod’, ‘Pot’ are not important in the diagnosis of CKD, rather they are important if any 1st line investigation provides useful analysis for further serum electrolyte analysis. Same goes for ‘Pcv’ which is used in much later stage prediction of CKD, not in initial diagnosis. Since ‘Hemo’ value is good enough for analyzing whether any patient has Anaemia or not, we do not separately need ‘rbc’ and ‘rbcc’. In stead of these features, our access to some features like Glomerular Filtration Rate (eGFR), Bi-Carbonate (HCO3), HbA1C (Mean Blood Glucose), Pro-BNP (B-type natriuretic peptide (BNP) test), S.Ph (Sodium Phosphate), PTH (Parathyroid Hormone) etc. would be more clinically significant variables. So, efforts were established for contacting the Apollo Hospital in Madurai, India several times, specially with the contributing doctor of the dataset but to no avail. So, the scope of this research, is limited to the choice of features by choosing the ones from the feature extraction techniques discussed before and discarding the ones suggested by Dr. Mahmood. So, the final 6 features for applying the Machine Learning models are:‘hemo’, ‘sc’, ‘al’, ‘htn’, ‘age’, ‘dm’. Moreover, we discussed with Dr. Mahmood about the missing data imputation process and she also agreed that the best way of imputation would be looking at similar patients data to get estimates about the features with missing values and using just any mathematical measures like mean, median, mode would not be clinically correct.
@@ IV Model Application Process
We applied SMOTE on the imputed CKD dataset so that both the ‘CKD’ and ‘Non-CKD’ instances were equal. After application of SMOTE on the dataset to tackle the class imbalance issue, Stratified K-Fold cross validation is applied to divide the dataset into equal K-folds for training and testing. Stratified K-Fold is performed to eradicate the random sampling issue for the train-test split. Also, it ensures that a different sample is always chosen for the test set and no sample gets ”missed out” while evaluation. For our analysis we chose k = 10, since we have 500 records after SMOTE application and with k = 10, k - 1 = 9 are the training set and 1 is the test set. So, 90% of 500 or 450 records would be training set for each fold and 50 records would be the test set for each record. With k = 10, after all the iterations, all the records in turn would act as test set which is previously unseen to it’s corresponding training set. So, the class imbalance issue is first of all tackled by using SMOTE and next by application of Stratified K-fold to get rid of random test case choosing. These methods ensure that stratified sampling is performed for model application and evaluation. The scope of this research lies fundamentally in prediciction of CKD first correctly given the independent set of variables and then application of XAI algorithms for explainability. For prediction task, the research problem is basically a binary classification task of predicting ”CKD” or ”Non-CKD”. Hence, we used Logistic Regression first and then Bayesian Estimation or Naive Bayes model with Stratified K-Fold for train-test split.
@@ V Model Performance on CKD Dataset
Table 7 and 8 indicate the scores of the applied models in the accuracy metrics. From the previous sections on feature selection through the 4 techniques in table 6 and discussion with expert Nephrologists, final 6 features were narrowed down to compare the prediction performance of the Machine Learning models in the whole feature set versus the selected subset of features. The models discussed in the model application process are evaluated on the selected 6 features.
Figure 2: Cumulative Confusion Matrix for Random Forest
We tested the model performance not just on the 4 accuracy metrics combining both classes, but also for the prediction of individual classes. We did not want to be in a scenario where the ML models performed very well in detection of one class and failed miserably in another still the accuracy scores looked really good. As evident from table 7, all the applied models have very high scores on the accuracy metrics. The number of true positives and true negatives are high and false positives and false negatives are also very low accounting for the high Precision-Recall score in all of the models. Overall, all the models have performed really well on the selected subset of 6 features for this dataset. This shows to indicate that the feature engineering process has yielded better selection of features than all combined 24 independent features. For the LR model, we chose to keep ‘L1’ regularization parameter to penalize the wrong results by using the absolute value of the co-effecients in the LR cost function. Also, we chose ”class weight” as ”balanced” since after performing SMOTE on the original dataset, we have a balanced positive and negative class in the dataset. For SVM model, grid search was performed to find the best parameters suited to this dataset. For the decision tree model, ”Gini Index” was used as an entropy measure to prune out branches from features since in the feature selection process we used entropy and variance to select best subset of features. Among all the models that are applied on both the entire dataset and selected subset of the features, Random Forest model has performed the best balancing all the 4 metric scores compared to all other models.
@@ V-A Random Forest Model Performance
Since of all the 7 models applied, Random Forest (RF) has performed the best in terms of accuracy metrics, this study explored more of this model performance in terms of confusion matrix, decision trees and then to the explainability.
For the Confusion matrix, we have selected to look at cumulative confusion matrix since we have used Stratified 10 fold cross validation for model evaluation. Since each fold had 50 samples in the test set, the cumulative number of samples was total 250. Looking at the confusion matrix for the Random Forest in figure 2, it is observed that 245 cases of True Positive (class 0 = CKD) and 248 cases of True Negative (class 1 = Non-CKD) prediction was found in the prediction set. On the contrary, 7 mis-classifications occurred with 2 False Positives and 5 False Negatives. Hence the Average Precision score is slightly lower than that of average Recall Score. So overall in total 250 cases of actual class of CKD, only 2 cases were there where the Random Forest model failed to correctly predict the ”CKD” class and identified the two subjects as ”Non-CKD” patients.
Figure 3 represents the Decision Trees generated for the Random Forests for the prediction model. The samples at the root or leaf nodes represent the percentage of samples that are reaching that node. The two values in the array represent the contribution of that feature in the prediction of Non-CKD and CKD. The root node at the top dignifies the feature value with the highest Gini Score and the next nodes at the leaves are the feature scores in terms of Gini Impurity in descending order of sample proportion. In our case, the root node starts from ‘al’ which is indicative of the strongest feature contributor. Then the decision tree branching occurs based on gini impurity or the entropy of each feature starting with ‘al’. If the ‘al’ value is less or equal to 0.587 only then the entropy value is 0.5 which is the highest for a binary classification task. At that value of ‘al’, the decision tree would branch out to the prediction process of the two classes. If the value of ‘al’ was higher than 0.587, the class prediction would be 0 or CKD right there. The next contributing feature is the ‘Hemo’ and it also branches out for the two class predictions when it’s value is less than or equal to 13.387. We can validate the peak value of ‘Hemo’ towards prediction of target class by analyzing figure 10 where it is observed that at 13.387, ‘Hemo’ is contributing most in prediction.
Figure 3: Decision Tree for Best Fold of Random Forest
Point to note here is that Random Forest model combines a number of Decision Trees and based on the best pruning results of the decision trees, the final results are presented in random forest. For this study, the Grid Search technique of the Random Forest model was also employed to find out best set of parameters for prediction. It was found that a depth of 14 with number of estimators as 472, the best Random Forest Model was selected. The best performing Random Forest model would be further evaluated using the feature importances in the training phase and the validation set would be explained using XAI tools.
During the training phase as depicted in figure 4, the Random Forest model has identified the “Hemo”,“Sc”, “Al” as the features with highest weights. The weights and coefficients of these features from the test and validation set with the help of XAI techniques were analyzed.
@@
In figure 5, a random prediction case was taken where the Random Forest model predicted that the patient has CKD. The strongly contributing features for the CKD confidence are ‘hemo’, ‘al’, ‘dm’, ‘sc’ etc. The range of values these features exhibit were also observed when the RF model predicts as CKD. For the CKD prediction in this testcase, ‘Hemoglobin’ values are less than 10.9. Presence of ‘Albumin’ is indicative of CKD which is represented by ‘al ¿ 2.0’ here. Presence of Diabetes indicated by ‘Dm’ value greater than 1 and high serum creatinine shown by ‘Sc’ ¿ 2.40.
For the Non-CKD prediction, LIME shows ‘al’, ‘sc’, ‘hemo’, ‘dm’, ‘htn’. Here, the LIME tool has visualized that for this instance of the patient, the best performing Random Forest model has identified it to be a Non-CKD patient because ‘Hemoglobin’ count in gms is greater than 13.5, low level of Albumin, less chances of Hypertension, Serum Creatinine in a very low range.
@@
Model is more confident about CKD prediction when “Sc” values are are greater than 2.40 and more confident about Non-CKD prediction when “Sc” values are less than 0.80. Lastly, the feature “Al” values are on the lower end for the non-CKD prediction and on the higher end for the CKD prediction as it should be as per clinical requirements.
Figure 7 shows the global impact of the features in the CKD prediction where the X-axis represents the impact score of the features and the Y-axis represents the feature values. It is observed that for high values of ‘hemo’ and ‘al’, generally they have high impact on CKD prediction. The impact of ‘age’ in CKD prediction is concentrated in a region. Features like ‘hemo’, ‘sc’, ‘al’ and ‘dm’ contribute more towards the prediction of CKD compared to ‘htn’ and ‘age’. Same kind of pattern is observed for Non-CKD prediction. Since SHAP provides us with the feature importances and values for the global test cases (Here we indicate the best test case chosen from 10 fold cross validation) unlike LIME, the two figures represent the feature importance scores and the feature values exhibited by the independent features for the contrasting class predictions.
Fig 8 depicts the scatter plots of global test cases to visualize the inter-dependence of independent features. Here, the top two features with respect to feature contribution i.e. “al” and “hemo” are analyzed. Presence of two clusters could be noticed in the inter-dependence plot where ‘hemo’ values range from 8 to 13 and ‘al’ values range from 1.7 and above. The 2nd cluster is for ‘hemo’ values ranging from 14 to 18 and ‘al’ values from 0 to 1.0. The two clusters indicate CKD vs Non-CKD prediction where the first cluster represents the CKD predictions for the test cases and the 2nd cluster represents the Non-CKD predictions.
@@
The peak values of prediction are also validated by the Accumulated Local Effects (ALE) plots in fig 10. In the X-axis, Haemoglobin level are plotted against the feature contribution scores in the Y-axis. It is seen that the feature contribution for ‘Hemo’ is the highest when the ‘hemo’ value is close to 13 and the contribution score remains constant with the increasing values.
@@
Figure 10: ALE Plot showing effect on prediction using ’Hemo’
Figure 11: PDP Plot showing effect on prediction using ’Hemo’ and ’Sc’
The Partial Dependence Plots (PDP) in figure 11 validate how the feature values dependency works in the predictions. If the partial dependence plot in fig 11 is analyzed, it would be seen that the more strongly contributing features ‘hemo’ has their peak contribution when ‘sc’ values are greater than 1.25 and ‘hemo’ values are greater than 15. ‘Hemo’ values also reach a local maxima at values 12.5 to 13. The combined dependence of the two features is also shown when local maxima is reached at 0.44 and 0.34 (in terms of contribution towards model prediction) when the ‘Hemo’ value is greater than 15, ‘Sc’ is greater than 1.25 and another at 0.55 when the ‘Sc’ is less than 1.25 and ‘Hemo’ value greater than 15.